The stock price for the Bare Bear Corp. (BBC) was $126.70 at the end of 2014, $121.50 at the end of 2015, $104.25 at the end of 2016, and $142.30 at the end of 2017. BBC paid dividends of $3.75 in 2015 and 2016, and $4.00 in 2017. What were the dollar returns and percentage returns on BBC’s stock for 2015, 2016 and 2017
Answer #1

dollar returns for 2015=121.50-126.70+3.75=-1.45

dollar returns for 2016=104.25-121.50+3.75=-13.50

dollar returns for 2017=142.30-104.25+4.00=42.05

percentage return for 2015=dollar returns for 2015/price as on 2014

=-1.45/126.70

=-1.14%

percentage return for 2016=dollar returns for 2016/price as on 2015

=-13.50/121.50

=-11.11%

percentage return for 2017=dollar returns for 2017/price as on 2016

=42.05/104.25

=40.34%

the above will be answer..